# InterCLIP-MEP: Interactive CLIP and Memory-Enhanced Predictor for Multi-modal Sarcasm Detection

### v5

@@ 3.1 Interactive CLIP
$\hat{y}=\textrm{softmax}(\mathcal{F}_{c}(\tilde{H}^{f})),\quad\hat{y}\in\mathbb{R}^{N\times 2},$ (5)
@@
$\mathcal{L}^{c}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i,1})+(1-y_{i})\log(1-\hat{y}_{i,0})\right],$
@@
$\hat{H}^{f}=\textrm{norm}(\mathcal{F}_{p}(\tilde{H}^{f})),\quad\hat{H}^{f}\in\mathbb{R}^{{N}\times d_{f}},$ (7)
@@
$\mathcal{L}^{p}=\textrm{mean}(\hat{H}^{f}_{P}\cdot\hat{H}^{f^{T}}_{N})+\textrm{mean}(1-\hat{H}^{f}_{P}\cdot\hat{H}^{f^{T}}_{P})\\ +\textrm{mean}(1-\hat{H}^{f}_{N}\cdot\hat{H}^{f^{T}}_{N}),$ (8)
@@
We fully train the modules $\mathcal{F}_{c}$, $\mathcal{F}_{p}$, the adapting projection layers ($\mathcal{F}_{t}$ and $\mathcal{F}_{v}$), the gated projection layers ($\mathcal{G}_{t}$ and $\mathcal{G}_{v}$), and the learnable gating parameters ($\beta_{t}$ and $\beta_{v}$). We use LoRA Hu et al. (2022) to fine-tune parts of the weight matrices $\mathbf{W}$ in the self-attention modules of all encoders, specifically various combinations of $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$. We consider $\mathbf{W}$ and the rank $r$ of LoRA as hyperparameters for our study. All learnable parts are optimized by minimizing the joint loss:
$\mathcal{L}=\mathcal{L}^{c}+\mathcal{L}^{p}.$ (9)
@@
Input: Memory size $L$, Learned InterCLIP model, classification module $\mathcal{F}_{c}$ and projection module $\mathcal{F}_{p}$
Output: Final prediction $\hat{y}^{p}$
1:  Initialize memory $\mathcal{M}\in\mathbf{0}^{2\times L\times d_{f}}$
2:  Initialize index $\mathcal{I}\in\mathbf{0}^{2}$
3:  Initialize entropy records $\mathcal{C}\in\mathbf{0}^{2\times L}$
4:  for $i\leftarrow 1$ to $N_{\textrm{test}}$ do
5:     $\tilde{h}^{f}_{i}\leftarrow\textrm{InterCLIP}(\mathcal{P}_{i})$
6:     $\hat{y}_{i}\leftarrow\textrm{softmax}(\mathcal{F}_{c}(\tilde{h}^{f}_{i}))$
7:     $\ell_{\textrm{pse}_{i}}\leftarrow\arg\max_{j}(\hat{y}_{i,j}),\;j\in\{0,1\}$
8:     $c_{i}\leftarrow-\hat{y}_{i,0}\log{\hat{y}_{i,0}}-\hat{y}_{i,1}\log{\hat{y}_{i,1}}$
9:     $\hat{h}^{f}_{i}\leftarrow\textrm{norm}(\mathcal{F}_{p}(\tilde{h}^{f}_{i}))$
10:     if $\mathcal{I}[\ell_{\textrm{pse}_{i}}]<L$ then
11:        $\mathcal{M}[\ell_{\textrm{pse}_{i}}][\mathcal{I}[\ell_{\textrm{pse}_{i}}]]\leftarrow\hat{h}^{f}_{i}$
12:        $\mathcal{C}[\ell_{\textrm{pse}_{i}}][\mathcal{I}[\ell_{\textrm{pse}_{i}}]]\leftarrow c_{i}$
13:        $\mathcal{I}[\ell_{\textrm{pse}_{i}}]\leftarrow\mathcal{I}[\ell_{\textrm{pse}_{i}}]+1$
14:     else
15:        $j\leftarrow\textrm{GetMaxIdx}(\mathcal{C}[\ell_{\textrm{pse}_{i}}])$
16:        if $c_{i}<\mathcal{C}[\ell_{\textrm{pse}_{i}}][j]$ then
17:           $\mathcal{M}[\ell_{\textrm{pse}_{i}}][j]\leftarrow\hat{h}^{f}_{i}$
18:           $\mathcal{C}[\ell_{\textrm{pse}_{i}}][j]\leftarrow c_{i}$
19:        end if
20:     end if
21:     $\textrm{logits}\leftarrow\left[\sum_{k=0}^{\mathcal{I}[0]}(\hat{h}^{f}_{i}\mathcal{M}[0]^{T})_{k},\sum_{k=0}^{\mathcal{I}[1]}(\hat{h}^{f}_{i}\mathcal{M}[1]^{T})_{k}\right]$
22:     $\hat{y}^{p}_{i}\leftarrow\textrm{softmax}(\textrm{logits})$
23:     yield $\hat{y}^{p}_{i}$
24:  end for
3.2 Memory-Enhanced Predictor
@@ 4 Experiment
Table 1: Statistics of MMSD and MMSD2.0.
Method MMSD2.0 MMSD
Acc. (%) F1 (%) P (%) R (%) Acc. (%) F1 (%) P (%) R (%)
Text
TextCNN Kim (2014) 71.61* 69.52* 64.62* 75.22* 80.03* 75.32* 74.29* 76.39*
Bi-LSTM Zhou et al. (2016) 72.48* 68.05* 68.02* 68.08* 81.90* 77.53* 76.66* 78.42*
SMSD Xiong et al. (2019) 73.56* 69.97* 68.45* 71.55* 80.90* 75.82* 76.46* 75.18*
RoBERTa Liu et al. (2019) 79.66* 76.21* 76.74* 75.70* 93.97* 92.45* 90.39* 94.59*
Image
ResNet He et al. (2016) 65.50* 57.58* 61.17* 54.39* 64.76* 61.53* 54.41* 70.80*
ViT Dosovitskiy et al. (2021) 72.02* 69.72* 65.26* 74.83* 67.83* 63.40* 57.93* 70.07*
Text-Image
HFM Cai et al. (2019) 70.57* 66.88* 64.84* 69.05* 83.44* 80.18* 76.57* 84.15*
Att-BERT Pan et al. (2020) 80.03* 77.04* 76.28* 77.82* 86.05* 82.92* 80.87* 85.08*
CMGCN Liang et al. (2022) 79.83* 76.90* 75.82* 78.01* 86.54* 84.09* -* -*
HKE Liu et al. (2022) 76.50* 72.25* 73.48* 71.07* 87.36* 72.25* 81.84* 86.48*
DIP Wen et al. (2023) 80.59 78.23 75.52 81.14 89.59 87.17 87.76 86.58
DynRT Tian et al. (2023) 70.37 68.55 63.02 75.15 93.59 91.93 90.30 93.62
Multi-view CLIP Qin et al. (2023) 85.64* 84.10* 80.33* 88.24* 88.33* 85.55* 82.66* 88.65*
G2SAM Wei et al. (2024) 79.43 78.07 72.04 85.20 90.48 88.48 87.95 89.02
InterCLIP-MEP (Ours)
InterCLIP-MEP w/ RoBERTa (Ours)
Table 2: Main results. We use * to indicate that the results are taken from Qin et al. (2023). - indicates that results are not reported. $\dagger$ indicates our reproduced results. Underlined values represent the best multi-modal baseline for comparison. Bold values indicate those that surpass the underlined baseline. $L_{2}$ for MMSD2.0 and $L_{1}$ for MMSD denote the optimal MEP memory sizes.
Multi-view CLIP  85.64  165  488  51  15.59
DIP  80.59  196  OOM  OOM  OOM
G2SAM  79.43  116  90  13  18.32
DynRT  70.37  25  370  26  8.03
InterCLIP-MEP  86.72  8  55  6  6.14
Table 3: Efficiency comparison of different methods. To demonstrate the efficiency of InterCLIP-MEP, we selected several recent baselines for comparison. The analysis was conducted using the MMSD2.0 dataset on a single NVIDIA RTX 4090 GPU with a batch size of 128. In the table, Fitting Time / Epoch indicates the time required for each epoch during training and validation and OOM indicates Out of Memory, referring to GPU memory overflow.
w/o Inter w/ TW w/ V2T w/ T2V
Variant Acc. (%) F1 (%) Acc. (%) F1 (%) Acc. (%) F1 (%) Acc. (%) F1 (%)
Baseline 86.05 84.81 85.51 84.26 86.26 85.00 86.72 85.61
w/o Proj 85.76 84.43 85.43 84.05 85.68 84.22 86.22 84.51
w/o MEP 85.39 83.99 85.22 83.79 86.26 84.78 86.26 84.82
w/o LoRA 82.44 77.73 76.42 74.37 73.31 72.22 75.13 71.79
Table 4: Ablation study of InterCLIP-MEP, with Baseline denoting results without ablation.
4.1 Experimental Settings
@@
Following Qin et al. (2023), we evaluate performance on MMSD Cai et al. (2019) and MMSD2.0 Qin et al. (2023) using accuracy (Acc.), precision (P), recall (R), and F1-score (F1) as metrics. We present the statistics of the two datasets in Table 1.
@@
We compare the effectiveness of the InterCLIP-MEP framework against several unimodal and multi-modal methods. For text modality methods, we compare with TextCNN Kim (2014), Bi-LSTM Zhou et al. (2016), SMSD Xiong et al. (2019), and RoBERTa Liu et al. (2019). For image modality methods, we compare with ResNet He et al. (2016) and ViT Dosovitskiy et al. (2021). We compare with state-of-the-art multi-modal methods, including HFM Cai et al. (2019), Att-BERT Pan et al. (2020), CMGCN Liang et al. (2022), HKE Liu et al. (2022), DIP Wen et al. (2023), DynRT Tian et al. (2023), Multi-view CLIP Qin et al. (2023), and G2SAM Wei et al. (2024), which employ various techniques such as hierarchical fusion, graph neural networks, and dynamic routing for multi-modal sarcasm detection.
4.2 Main Results
To validate the effectiveness of our InterCLIP-MEP framework, we conduct experiments using the original CLIP as the backbone instead of InterCLIP, referred to as w/o Inter. We compare this configuration with three interaction modes of InterCLIP: w/ V2T, w/ T2V, and w/ TW.
For each experiment, we condition only the top four layers of the self-attention modules, with the projection dimension $d_{f}$ set to 1024. We set the LoRA rank $r$ to 8, fine-tuning the self-attention weight matrices $\mathbf{W}$, specifically $W_{k}$, $W_{v}$, and $W_{o}$. For the memory size $L$, we select the optimal size from a fixed set of candidate values $\mathbf{L}$  [^2] The fixed candidate values are {128, 256, 384, 512, 640, 768, 896, 1024, 1152, 1280}.. The main results are shown in Table 2.
Performance on MMSD2.0.
For MMSD2.0, our framework consistently outperforms or matches the performance of state-of-the-art methods, whether using InterCLIP or the original CLIP as the backbone, as shown in Table 2 (InterCLIP-MEP). This demonstrates the effectiveness of our training strategy and MEP. Our results show that w/ V2T and w/ T2V outperform w/o Inter, demonstrating that InterCLIP captures text-image interactions more effectively. Furthermore, w/ T2V achieves superior performance compared to w/ V2T, likely due to the inherent complexity of the visual space, which presents challenges for the projection layer when mapping vision representations into the text encoder space. In contrast, w/ TW performs worse than other configurations, possibly because embedding representations within both encoders increases the learning difficulty. In summary, InterCLIP with T2V interaction, combined with our training strategy and MEP, delivers the most promising results. These findings underscore the robustness and adaptability of our framework, establishing it as a highly effective solution for capturing nuanced text-image interactions and addressing the complexities of multi-modal sarcasm detection.
Performance on MMSD.
For MMSD, the RoBERTa-based text modality baseline significantly outperforms other methods due to spurious cues in the text, enabling accurate predictions solely based on textual features Qin et al. (2023). Consequently, models like DynRT, G2SAM, and DIP, which utilize RoBERTa or BERT for text feature extraction, achieve high performance on MMSD but experience a significant drop on MMSD2.0. To further investigate, we conduct an additional experiment, InterCLIP-MEP w/ RoBERTa, replacing the original text encoder with RoBERTa. While this change led to state-of-the-art performance on MMSD, it disrupted InterCLIP’s modality alignment, causing a reasonable performance drop on MMSD2.0. This suggests that MMSD’s text data contains spurious cues that allow models to rely heavily on the text encoder, while MMSD2.0, having been cleaned, requires more robust multi-modal capabilities. We further find that the w/ V2T variant consistently outperforms others in both InterCLIP-MEP and InterCLIP-MEP w/ RoBERTa experiments, underscoring the model’s tendency to overly depend on text modality.
Efficiency Comparison.
Our training strategy demonstrates both remarkable effectiveness and outstanding efficiency. To validate this, we conducted a comprehensive comparative analysis against leading state-of-the-art methods, as detailed in Table 3. For instance, the Multi-view CLIP method Qin et al. (2023) employs a multi-layer Transformer encoder for feature fusion, which, while effective, introduces a significant number of trainable parameters. This results in slower training and inference speeds and greater memory consumption. Similarly, the DIP method Wen et al. (2023) caches historical samples during training, which hinders its ability to support large-batch training under limited resource conditions. In contrast, our method operates with a batch size of 128 while utilizing a significantly smaller number of trainable parameters, which translates to notably faster training and validation cycles. Furthermore, by incorporating minimal parameter modifications to adapt CLIP and utilizing simple yet effective linear layers for representation fusion, our approach achieves superior inference speeds and drastically reduced memory consumption. These results highlight the practicality of our framework, establishing it as a benchmark for both computational efficiency and performance in multi-modal sarcasm detection.
4.3 Analysis of InterCLIP-MEP
Ablation study.
We remove the projection module $\mathcal{F}_{p}$ and train only the classification module $\mathcal{F}_{c}$ for prediction, denoted as w/o Proj. To test the necessity of using LoRA Hu et al. (2022) for fine-tuning, we keep the rest of InterCLIP-MEP unchanged and freeze all self-attention weights of InterCLIP, denoted as w/o LoRA, always selecting the optimal memory size $L$ for the MEP during inference. To evaluate the effectiveness of the MEP, we train both $\mathcal{F}_{p}$ and $\mathcal{F}_{c}$ but use only $\mathcal{F}_{c}$ during inference, denoted as w/o MEP.
Table 4 reports all results. All variants show performance declines compared to the baseline, demonstrating the importance of each module in InterCLIP-MEP. For w/ TW and w/o Inter, the w/o MEP variant performed worse than the w/o Proj variant. However, for w/ T2V and w/ V2T, the w/o MEP variant performs better than the w/o Proj variant. This suggests that backbones with strong image-text interaction capabilities benefit from training the classification module $\mathcal{F}_{c}$ along with the projection module $\mathcal{F}_{p}$, even without using MEP during inference. We also find that not using LoRA to fine-tune the self-attention modules results in significant performance loss, indicating that the original CLIP’s vanilla space is not directly suitable for the sarcasm detection task.
Case study.
As shown in Figure 4, we select three examples to further demonstrate the robustness of InterCLIP-MEP. We observe that direct predictions through the classification module $\mathcal{F}_{c}$ result in high prediction entropy and incorrect outcomes. However, MEP effectively mitigates the issue for cases where $\mathcal{F}_{c}$ fails to correctly identify the results by using the historical knowledge of the test samples. This integration ensures that even in complex situations, the model maintains a high level of accuracy.
Figure 4: Case study of InterCLIP-MEP. In the figure, GT represents the labels annotated by human experts, MEP represents the labels predicted by the memory-enhanced predictor, and $\mathcal{F}_{c}$ represents the labels provided by the classification module.
Figure 5: Visual comparison between InterCLIP and original CLIP. Both InterCLIP and the original CLIP were fine-tuned using the same training set and identical parameters. The key distinction is that the original CLIP does not incorporate interaction.
Visulization.
To further validate that InterCLIP is capable of more effectively capturing the interactive information between text and images compared to original CLIP, thereby aiding in the detection of sarcasm cues, we use GradCAM Selvaraju et al. (2017) to visualize the areas of focus during the inference process of the visual model in Figure 5. We observe that in the first example, which complains about the unpleasant odor caused by the sea lions, InterCLIP focuses more accurately on the location of the sea lions compared to the original CLIP. In the second example, which complains about traffic congestion, InterCLIP correctly focuses on the distribution of cars on the road, whereas the original CLIP’s focus is scattered.
5 Conclusion
In this paper, we propose InterCLIP-MEP, a novel framework for multi-modal sarcasm detection that directly addresses the challenges of modeling nuanced text-image interactions and managing prediction uncertainty. We design Interactive CLIP (InterCLIP) to embed cross-modal information within text and image encoders, enabling a deeper understanding of sarcasm cues. Additionally, we develop a Memory-Enhanced Predictor (MEP) to dynamically leverage historical sample knowledge, making our inference process more robust and adaptive. Through extensive experiments on MMSD and MMSD2.0 benchmarks, we demonstrate that InterCLIP-MEP achieves state-of-the-art performance while significantly reducing computational costs. By requiring fewer trainable parameters and less GPU memory, our method offers a lightweight, efficient, and scalable solution, setting a new benchmark for multi-modal sarcasm detection.
Limitations
While InterCLIP-MEP delivers strong performance, there remain areas for further refinement. For instance, the framework could benefit from additional techniques to better capture sarcasm that heavily relies on subtle cultural or highly specific contextual cues. Moreover, extending its application to more diverse and less structured datasets could be an interesting direction for future work, further broadening its practical applicability.
Ethical Considerations
This work focuses on advancing multi-modal sarcasm detection to improve understanding of complex communication in online content. While the proposed framework enhances detection accuracy and efficiency, potential misuse must be considered. Automated sarcasm detection could inadvertently amplify biases present in training data or be deployed for unethical purposes, such as targeted content moderation or surveillance. To mitigate these risks, we encourage the responsible use of this technology and emphasize the importance of using diverse and unbiased datasets during training to minimize unintended consequences. Furthermore, this research strictly adheres to ethical guidelines for data collection and usage.
Appendix A Implementation Details
@@
The model training and testing were conducted using PyTorch Lightning [^3] https://lightning.ai/. InterCLIP was constructed by leveraging the Transformers library Wolf et al. (2020). For the MMSD2.0 experiments, the initial weights for InterCLIP are based on clip-vit-base-patch32444https://huggingface.co/openai/clip-vit-base-patch32. For the MMSD experiments, we utilized the roberta-ViT-B-32 model architecture provided by OpenCLIP [^5] https://github.com/mlfoundations/open_clip, with the pre-trained checkpoint laion2b_s12b_b32k [^6] https://huggingface.co/laion/CLIP-ViT-B-32-roberta-base-laion2B-s12B-b32k. Custom scripts were developed to adapt its format to the Transformers library, ensuring compatibility with our framework. The model parameters were optimized using AdamW Loshchilov and Hutter (2019), with a learning rate set to 1e-4 for the LoRA fine-tuning modules and 5e-4 for other trainable modules. A cosine annealing scheduler with warmup was employed to dynamically adjust the learning rate, where the warmup steps constituted the first 20% of the total optimization steps, and the minimum learning rate was set to 1% of the initial rate. For the modules $\mathcal{G}_{t/v}$, $\mathcal{F}_{t/v}$, $\mathcal{F}_{c}$, and $\mathcal{F}_{p}$, simple multi-layer perceptrons (MLPs) were utilized. The training processing was performed with a batch size of 64 for 3 epochs. All experiments were run on a machine equipped with an NVIDIA RTX 4090 GPU.
Appendix B Experimental Details
B.1 Hyperparameter details
We summarize the hyperparameters involved in InterCLIP-MEP and their descriptions in Table 5. The hyperparameter settings for obtaining the main results in the paper are summarized in Table 6. For other baseline methods, we follow the optimal hyperparameter settings they reported.
@@ B.1 Hyperparameter details
Table 5: Summary of hyperparameters.
Param. Value

### v7

@@ 3.2. Training Strategy
(7) $\hat{y}=\textrm{softmax}(\mathcal{F}_{c}(\tilde{H}^{f})),\quad\hat{y}\in\mathbb{R}^{N\times 2},$
@@
(8) $\mathcal{L}^{c}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i,1})+(1-y_{i})\log(1-\hat{y}_{i,0})\right],$
@@
(9) $\hat{H}^{f}=\textrm{norm}(\mathcal{F}_{p}(\tilde{H}^{f})),\quad\hat{H}^{f}\in\mathbb{R}^{{N}\times d_{f}},$
@@
(10) $\mathcal{L}^{p}=\textrm{mean}(\hat{H}^{f}_{P}\cdot\hat{H}^{f^{T}}_{N})+\textrm{mean}(1-\hat{H}^{f}_{P}\cdot\hat{H}^{f^{T}}_{P})+\textrm{mean}(1-\hat{H}^{f}_{N}\cdot\hat{H}^{f^{T}}_{N}),$
@@
We fully train the modules $\mathcal{F}_{c}$, $\mathcal{F}_{p}$, the adapting projection layers ($\mathcal{F}_{t}$ and $\mathcal{F}_{v}$), the gated projection layers ($\mathcal{G}_{t}$ and $\mathcal{G}_{v}$), and the learnable gating parameters ($\beta_{t}$ and $\beta_{v}$). We use LoRA (edward2022lora) to fine-tune parts of the weight matrices $\mathbf{W}$ in the self-attention modules of all encoders, specifically various combinations of $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$. We consider $\mathbf{W}$ and the rank $r$ of LoRA as hyperparameters for our study. All learnable parts are optimized by minimizing the joint loss:
(11) $\mathcal{L}=\mathcal{L}^{c}+\mathcal{L}^{p}.$
3.3. Memory-Enhanced Predictor
@@
1:Input: Memory size $L$, Learned InterCLIP model, classification module $\mathcal{F}_{c}$ and projection module $\mathcal{F}_{p}$
2:Output: Final prediction $\hat{y}^{p}$
3: Initialize memory $\mathcal{M}\in\mathbf{0}^{2\times L\times d_{f}}$
4: Initialize index $\mathcal{I}\in\mathbf{0}^{2}$
5: Initialize entropy records $\mathcal{C}\in\mathbf{0}^{2\times L}$
6:for $i\leftarrow 1$ to $N_{\textrm{test}}$ do
7:  $\tilde{h}^{f}_{i}\leftarrow\textrm{InterCLIP}(\mathcal{P}_{i})$
8:  $\hat{y}_{i}\leftarrow\textrm{softmax}(\mathcal{F}_{c}(\tilde{h}^{f}_{i}))$
9:  $\ell_{\textrm{pse}_{i}}\leftarrow\arg\max_{j}(\hat{y}_{i,j}),\;j\in\{0,1\}$
10:  $c_{i}\leftarrow-\hat{y}_{i,0}\log{\hat{y}_{i,0}}-\hat{y}_{i,1}\log{\hat{y}_{i,1}}$
11:  $\hat{h}^{f}_{i}\leftarrow\textrm{norm}(\mathcal{F}_{p}(\tilde{h}^{f}_{i}))$
12:  if $\mathcal{I}[\ell_{\textrm{pse}_{i}}]<L$ then
13:   $\mathcal{M}[\ell_{\textrm{pse}_{i}}][\mathcal{I}[\ell_{\textrm{pse}_{i}}]]\leftarrow\hat{h}^{f}_{i}$
14:   $\mathcal{C}[\ell_{\textrm{pse}_{i}}][\mathcal{I}[\ell_{\textrm{pse}_{i}}]]\leftarrow c_{i}$
15:   $\mathcal{I}[\ell_{\textrm{pse}_{i}}]\leftarrow\mathcal{I}[\ell_{\textrm{pse}_{i}}]+1$
16:  else
17:   $j\leftarrow\textrm{GetMaxIdx}(\mathcal{C}[\ell_{\textrm{pse}_{i}}])$
18:   if $c_{i}<\mathcal{C}[\ell_{\textrm{pse}_{i}}][j]$ then
19:    $\mathcal{M}[\ell_{\textrm{pse}_{i}}][j]\leftarrow\hat{h}^{f}_{i}$
20:    $\mathcal{C}[\ell_{\textrm{pse}_{i}}][j]\leftarrow c_{i}$
21:   end if
22:  end if
23:  $\textrm{logits}\leftarrow\left[\sum_{k=0}^{\mathcal{I}[0]}(\hat{h}^{f}_{i}\mathcal{M}[0]^{T})_{k},\sum_{k=0}^{\mathcal{I}[1]}(\hat{h}^{f}_{i}\mathcal{M}[1]^{T})_{k}\right]$
24:  $\hat{y}^{p}_{i}\leftarrow\textrm{softmax}(\textrm{logits})$
25:  yield $\hat{y}^{p}_{i}$
26:end for
@@
4. Experiments
4.1. Experimental Settings
@@
Following (qin2023mmsd2), we evaluate performance on MMSD (Cai2019MultiModalSD) and MMSD2.0 (qin2023mmsd2) using accuracy (Acc.), precision (P), recall (R), and F1-score (F1) as metrics. We present the statistics of the two datasets in Table 2.
@@
We compare the effectiveness of the InterCLIP-MEP framework against several unimodal and multi-modal methods. For text modality methods, we compare with TextCNN (kim-2014-convolutional), Bi-LSTM (zhou-etal-2016-attention), SMSD (xiong2019sarcasm), and RoBERTa (liu2019roberta). For image modality methods, we compare with ResNet (he2015deep) and ViT (dosovitskiy2020image). We compare with state-of-the-art multi-modal methods, including HFM (Cai2019MultiModalSD), Att-BERT (pan-etal-2020-modeling), CMGCN (Liang2022MultiModalSD), HKE (Liu2022TowardsMS), DIP (wen2023dip), DynRT (Tian2023DynamicRT), Multi-view CLIP (qin2023mmsd2), and G2SAM (wei2024g2sam), which employ various techniques such as hierarchical fusion, graph neural networks, and dynamic routing for multi-modal sarcasm detection.
4.2. Implementation Details
@@
The model training and testing were conducted using PyTorch Lightning [^1] https://lightning.ai/. InterCLIP was constructed by leveraging the Transformers library (wolf-etal-2020-transformers). For the MMSD2.0 experiments, the initial weights for InterCLIP are based on clip-vit-base-patch32222https://hf.co/openai/clip-vit-base-patch32. For the MMSD experiments, we utilized the roberta-ViT-B-32 model architecture provided by OpenCLIP [^3] https://github.com/mlfoundations/open_clip, with the pre-trained checkpoint laion2b_s12b_b32k [^4] https://hf.co/laion/CLIP-ViT-B-32-roberta-base-laion2B-s12B-b32k. Custom scripts were developed to adapt its format to the Transformers library, ensuring compatibility with our framework. The model parameters were optimized using AdamW (loshchilov2017decoupled), with a learning rate set to 1e-4 for the LoRA fine-tuning modules and 5e-4 for other trainable modules. A cosine annealing scheduler with warmup was employed to dynamically adjust the learning rate, where the warmup steps constituted the first 20% of the total optimization steps, and the minimum learning rate was set to 1% of the initial rate. For the modules $\mathcal{G}_{t/v}$, $\mathcal{F}_{t/v}$, $\mathcal{F}_{c}$, and $\mathcal{F}_{p}$, simple multi-layer perceptrons (MLPs) were utilized. The training processing was performed with a batch size of 64 for 3 epochs. All experiments were run on a machine equipped with an NVIDIA RTX 4090 GPU.
Table 3. Summary of hyperparameters.
@@
Table 4. Hyperparameter settings.
Parameter Value